# OpenFACADES: An Open Framework for Architectural Caption and Attribute Data Enrichment via Street View Imagery

### v1

@@ 3.3 Establishing dataset and multimodal models
Street-level building dataset
@@
Following the previous process, building information is assigned to the detected buildings in the imagery. In this study, we focus on the classification of building type, facade material, construction year (age), and number of floors, which have been identified as essential attributes in prior studies and are supported by relatively sufficient data for model training and evaluation. Specifically, we utilize labels from building data corresponding to the categories: building, building:material, start_date, and building:levels, which are primarily sourced from OSM and supplemented by additional building datasets, as mentioned in Section 3.1. Here, building type and facade material are treated as categorical variables, while construction year and number of floors are represented as numerical values.
@@
From the full set of building data, we sample buildings with available category labels to construct the dataset for subsequent model development. The dataset is assembled and divided into training and test sets based on the following three principles: (1) ensuring sufficient labels across all classes to avoid biased predictive accuracy in machine learning tasks, and balanced class distribution in test set; (2) maintaining a balanced geospatial distribution across cities to represent the diversity of architectural designs; and (3) preventing the same building from appearing in both the training and test sets to minimize data leakage.
@@
The dataset contains four types of labels: categorical, single-word Q&A, multi-attribute Q&A, and captioning labels. Categorical labels are used to fine-tune CNN/ViT-based vision models, serving as the baseline for evaluating the performance of common practices. Single-word Q&A labels are derived from categorical labels by appending the label to specific questions about the four building attributes, thereby generating concise question-to-label pairs based on building information. Multi-attribute Q&A and captioning labels are generated using the state-of-the-art multimodal large language model, ChatGPT-4o, through the OpenAI API [^3] https://www.openai.com/. This task involves prompting the model to annotate or describe the building features visible in the images, thereby creating an image-text training set. Table 2 provides detailed indication of data sources and examples of these labels, showcasing the comprehensive structure of the dataset designed to facilitate the interpretation of building imagery.
@@
Table 2: Different label types and data collection approaches for developing a street-level building dataset.
Label type Target model Data source Examples
Categorical label Computer vision models OpenStreetMap + government data apartment, house, retail, etc.
Single-word Q&A Vision-Language models OpenStreetMap + government data Q: Assign a brief label for the floor count of the building in the image. A: 12
Multi-attribute Q&A Teacher model (ChatGPT-4o) Q: Provide concise labels for each category using the following JSON format. Select appropriate values from the provided options for each category: {“building_type”: “(choose one option from: ‘apartments’, ‘house’, ‘retail’, ‘office’, ‘hotel’, ‘industrial’, ‘religious’, ‘education’, ‘public’, ‘garage’)”, “alternate_building_type”: “(choose one option from: ‘apartments’, ‘house’, ‘retail’, ‘office’, ‘hotel’, ‘industrial’, ‘religious’, ‘education’, ‘public’, ‘garage’)”, “building_age”: “(a 4-digit year indicating the approximate construction date of the building)”, “floors”: “(a numeric value representing the total number of floors)”, “surface_material”: “(choose one option from: ‘brick’, ‘wood’, ‘concrete’, ‘metal’, ‘stone’, ‘glass’, ‘plaster’)”, “alternate_surface_material”: “(choose one option from: ‘brick’, ‘wood’, ‘concrete’, ‘metal’, ‘stone’, ‘glass’, ‘plaster’)”} A: {“building_type”: “office”, “alternate_building_type”: “retail”, “building_age”: 1950, “floors”: 8, “surface_material”: “brick”, “alternate_surface_material”: “stone”}
Captioning Teacher model (ChatGPT-4o) Q: Analyze the building shown in the image and provide a detailed description of its architectural features. Then, describe the building type, the building’s age (by specifying an approximate construction year), the primary facade material (the main material visible on the building’s surface), the construction material, and the total number of floors in the building. A: The building in the image appears to be a residential structure, likely a house …… In summary, the building is a traditional wooden house with a steeply pitched roof, likely constructed in the late 19th to early 20th century. It has a single-story design with a wooden facade and is well-maintained.
Vision-language models
@@
To address the limitations of traditional categorical classification models in building attribute analysis, we leverage InternVL2.5 (Chen et al., 2024b), an open-source multimodal large language model (MLLM) designed for unified visual-language reasoning. As depicted in Figure 5, InternVL2.5 is built on the “ViT-MLP-LLM” paradigm by integrating a scalable vision encoder (InternViT) (Chen et al., 2024c), a multi-layer perceptron (MLP) projector, and a large language model (LLM). The vision encoder is InternViT-300M-448px-V2.5, a distilled variant of the 6B-parameter model optimized via dynamic high-resolution training and next token prediction (NTP) loss (Chen et al., 2024b). This architecture processes 448×448 pixel image tiles through a pixel unshuffle operation, reducing 1024 visual tokens to 256 for efficient cross-modal alignment.
@@
The model is selected for its general-purpose captioning and open-vocabulary classification capabilities, critical for capturing the multifaceted attributes of buildings (e.g., material, style, type) within a unified framework. Unlike conventional models restricted to predefined labels, InternVL2.5’s contrastive vision-language pretraining enables semantic reasoning over diverse facade characteristics, aligning with our goal of holistic building profiling. Full-model tuning is conducted through optimizing three components (Figure 5): (1) InternViT-300M Vision Encoder: Retrained on street-level building images to enhance facade feature extraction, leveraging dynamic high-resolution (448px) inputs; (2) MLP Projector: Adjusted to align building-specific visual tokens with textual embeddings in the LLM space; (3) LLM Head: Fine-tuned using the corpus of building characteristic descriptions to generate structured captions.
@@
After fine-tuning, we design a comprehensive set of experiments to evaluate the model’s performance in several respects: its general capability of VLMs, its sensitivity to training data size (ablation analysis), and its effectiveness in inferring building attributes across different cities and categories. To align these investigations more closely with real-world building profiling practices, we further conduct comparative assessments involving both VLMs and conventional categorical classification models. These comparisons target three main objectives: (1) examining the performance of the fine-tuned models, (2) assessing their generalizability to unseen data (a crucial concern in cross-city studies, where models trained on one urban area may not perform consistently when applied elsewhere (Sun et al., 2022b; Hou et al., 2025)), and (3) evaluating robustness to heterogeneous noises and degradation. The latter is particularly relevant because crowdsourced images, unlike standardized remote sensing imagery, frequently exhibit diverse quality issues (Hou and Biljecki, 2022). These experiments offer deeper insights into models’ generalization capacity.
@@
To investigate robustness and impact of common issues in imagery, we adopt the methodology outlined by Hendrycks and Dietterich (2019), which measures model resilience against common image corruptions and perturbations. Guided by the image quality criteria proposed by Hou and Biljecki (2022), we algorithmically generate four types of corruptions (Figure 6) — occlusion, motion blur, Gaussian noise, and brightness alterations — and apply them to the test set. We then evaluate each model’s performance under these degraded conditions using Relative Corruption Errors ($Relative$ CE) (Hendrycks and Dietterich, 2019). First, the baseline error rate $E^{m}_{clean}$ was determined for model $m$ on the uncorrupted data. Next, we compute the error rate $E^{m}_{c,s}$ for each corruption type $c$ at severity level $s$ (1 $\leq$ s $\leq$ 3). In classification tasks (building type and surface material), the error rate is defined as $1-Accuracy$, whereas for regression tasks (predicting number of floors and building age), it is defined as $1-R^{2}$. Finally, to account for the varying difficulties introduced by each corruption, we normalize these error rates by dividing by the ResNet50 baseline error; $Relative$ CE is calculated as:
@@
This normalization provides a clearer measure of how much each model’s performance declines under different corruptions. Averaging these $Relative$ CE from four types of corruptions results in the $Relative$ $m$CE, which represents the overall relative performance degradation when the models encountering corruptions.
@@
4 Experiments and results
@@ 4.1 Street-level building dataset
The building dataset is established using panoramic images sourced from Mapillary [^4] https://www.mapillary.com. We manually select cities that have a sufficient number of panoramic images available through the Mapillary online interface, and that also have a considerable amount of objective building attributes openly in OSM. Ultimately, seven cities from three continents are chosen, including Amsterdam, Berlin, Helsinki, San Francisco, Washington D.C., Houston and Manila, balancing the dataset across both selection aspects. Among them, Helsinki is selected due to its rich availability of building material data from the Buildings in Helsinki data [^5] https://hri.fi/data/en_GB/dataset/helsingin-rakennukset, while Amsterdam provides diverse data on building age, to add sufficient data on according aspects.
The metadata for panoramic SVIs is first downloaded within the defined city boundaries using the Mapillary Python Software Development Kit [^6] https://github.com/mapillary/mapillary-python-sdk, while building data is retrieved using OSMnx (Boeing, 2017). Subsequently, the data undergoes the process described in Section 3.1 to calculate the angle of view, evaluate observation quality, and identify candidate images. These selected images are then utilized for building detection, image reprojection, and filtering, as detailed in Section 3.2, resulting in a collection of individual building images for each city. Table 3 provides a detailed breakdown of the number of buildings, SVI images retrieved, individual buildings detected with associated images, and the ratio of completeness for each city. While completeness varies among cities due to differences in the availability and quality of Mapillary images uploaded for specific locations, around 50% of buildings in city centers can be observed and analyzed.
@@
City Total building footprints Total images retrieved Total individual building images Buildings with images Percentage detected City center completeness (2.5km×$\times$2.5km)
@@
As discussed in Section 3.3, building images with available attributes are sampled to construct a class-sufficient dataset for model development, resulting in a total of 30,180 images. Figure 7 illustrates the distribution of images across relevant categories for each attribute, comprising 17,530 images for building type, 2,871 for surface material, 7,228 for floors, and 5,927 for age. The dataset is divided into training, validation, and test sets in a 6:1:3 ratio. For the training and test sets, ChatGPT-4o is employed to generate multi-attribute Q&A and captioning labels, enriching the dataset with additional descriptive annotations for both training and benchmarking. To comprehensively capture ambiguous architectural features, we also obtain an alternative classifier by using the prompts “alternate_building_type” and “alternate_surface_material”, which represent the top-two predictions of the MLLMs.
@@
4.2 Vision-language model evaluation

### v2

@@ 3.3 Establishing dataset and multimodal models
Street-level building dataset.
@@
Following the previous process, building information is assigned to the detected buildings in the imagery. In this study, we focus on building type, facade material, construction year (age), and number of floors, which have been identified as essential attributes in prior studies and are supported by relatively sufficient data for model training and evaluation. Specifically, we utilize labels from building data corresponding to the categories: building, building:material, start_date, and building:levels in OSM alongside supplementary datasets, as mentioned in Section 3.1. Here, building type and facade material are treated as categorical variables, while construction year and number of floors are represented as numerical values.
@@
From the full set of building data, we sample buildings with available category labels to construct the dataset for subsequent model development. The dataset is assembled and divided into training, validation, and test sets based on the following three principles: (1) ensuring sufficient labels across all classes to avoid biased predictive accuracy; (2) maintaining a balanced geospatial distribution across cities to represent the diversity of architectural designs; and (3) preventing the same building from appearing in both different sets to minimize data leakage.
@@
The dataset contains four types of labels: single-attribute label, single-attribute Q&A, multi-attribute Q&A, and captioning labels. Single-attribute labels are used to fine-tune CV models, serving as the baseline for evaluating the performance of common practices. Single-attribute Q&A labels are derived from those single-attribute labels by appending the label to specific questions about the four building attributes, thereby generating concise question-to-label pairs based on building information. Multi-attribute Q&A and captioning labels are generated using the state-of-the-art multimodal large language model, ChatGPT-4o , through the OpenAI API [^3] https://www.openai.com/. This task involves prompting the model to annotate or describe the building features visible in the images, thereby creating an image-text dataset. Figure 5 provides detailed indication of data sources and examples of these labels.
Among these labels, the single- and multi-attribute Q&A outputs share a consistent, structured format constrained by predefined vocabulary, whereas the captioning task provides a free-form textual description that often embeds the same attributes in more expressive language, forming a hierarchical relationship with the Q&A formats. For example, a building labeled as “brick” (material), “1920” (age), “3” (floors), and “house” (type) in the Q&A tasks might be described in the captioning output as “a three-story house with visible brickwork, built in the early 20th century”, reflecting increased richness in narrative form. Inconsistencies can arise between OSM ground truth labels and ChatGPT-generated labels that slight discrepancies (e.g., differing building floor or age estimates) occur when ChatGPT assigns different labels to buildings with ambiguous features.
@@
Figure 5: Different label types and data collection approaches for developing a street-level building dataset. Data: (c) OSM and Mapillary contributors.
Vision-language models.
@@
To address the limitations of traditional CV models in building attribute analysis, we leverage InternVL3 (Zhu et al., 2025), an open-source MLLM designed for unified visual-language reasoning. As depicted in Figure 6, InternVL3 is built on the “ViT-MLP-LLM” paradigm by integrating a scalable vision encoder (InternViT) (Chen et al., 2024c), a multi-layer perceptron (MLP) projector, and a large language model (LLM). The vision encoder is InternViT-300M-448px-V2.5, a distilled variant of the 6B-parameter model optimized via dynamic high-resolution training and next token prediction (NTP) loss (Chen et al., 2024b). This architecture processes 448×448 pixel image tiles through a pixel unshuffle operation, reducing 1024 visual tokens to 256 for efficient cross-modal alignment.
@@
The model is selected for its general-purpose captioning and open-vocabulary classification capabilities, critical for capturing the multifaceted attributes of buildings (e.g., material, style, type) within a unified framework. Unlike conventional models restricted to predefined labels, InternVL’s contrastive vision-language pretraining enables semantic reasoning over diverse facade characteristics, aligning with our goal of holistic building profiling. Full-model tuning is conducted through optimizing three components (Figure 6): (1) InternViT-300M Vision Encoder: Retrained on street-level building images to enhance facade feature extraction, leveraging dynamic high-resolution (448px) inputs; (2) MLP Projector: Adjusted to align building-specific visual tokens with textual embeddings in the LLM space; (3) LLM Head: Fine-tuned using the corpus of building characteristic descriptions to generate structured captions. Here, we recast façade profiling as a multi-attribute text prediction task: all attributes and captions are encoded in a fixed-template prompt, and the MLLM is fine-tuned to generate that structured output. A single token-level cross-entropy loss implicitly handles both attribute prediction and free-form captioning, without requiring custom loss functions.
@@
4 Experimental settings
4.1 Implementations
To implement methods mentioned in Section 3.1, we manually select cities that have a sufficient number of panoramic images available through the Mapillary online interface, and that also have a considerable amount of objective building attributes openly. Ultimately, seven cities from three continents are chosen, including Amsterdam, Berlin, Helsinki, San Francisco, Washington D.C., Houston and Manila, balancing the dataset across both selection aspects. Among them, Helsinki is selected due to its rich availability of building material data from the Buildings in Helsinki data [^4] https://hri.fi/data/en_GB/dataset/helsingin-rakennukset, while Amsterdam provides diverse data on building age, to add sufficient data on according aspects.
The metadata for panoramic SVIs is first downloaded within the defined city boundaries using the Mapillary Python Software Development Kit [^5] https://github.com/mapillary/mapillary-python-sdk, while building data is retrieved using OSMnx (Boeing, 2017) and Overture Maps. Subsequently, the data undergoes the process described in Section 3.1 to calculate the angle of view, evaluate observation quality, and identify candidate images. These selected images are then utilized for building detection, image reprojection, and filtering, as detailed in Section 3.2, resulting in a collection of individual building images for each city.
To ensure the collection of high-quality building images, the filtering criteria we applied are detailed in Table 2. We acknowledge that stringent filtering may inadvertently exclude images from certain regions, e.g., areas where dense greenery obstructs building facades. To address this and support broader applicability, we have made our code flexible, allowing users to adjust filtering thresholds to accommodate different urban contexts and mitigate potential geographic bias. Additionally, other image features, such as distance between the building and the vantage point, visibility coverage of the building perimeter, and other indicators derived from semantic segmentation and scene classification, are generated during the processing pipeline. These features can be further integrated into filtering criteria in future implementation.
Table 2: Image filtering criteria used for selecting building images in this study.
Feature Description Selection Rule
Mapillary metadata
Image Type Select only panoramic images is_pano = True
Location (Duplicate) Remove images taken at the same location Filter duplicates based on coordinates
Image Quality Use Mapillary quality score $\in$ [0, 1] to retain images with valid or no score quality_score $\geq$ 0.2 or
quality_score = 0
Building metadata
Building Area Retain buildings with sufficient spatial footprint area >> 20 m2
Underground Exclude buildings fully underground building:levels $\geq$ 0 (OSM) or
is_underground = False (Overture)
Isovist analysis
Angle of View (AOV) Exclude extreme observation angles 10° $\leq$ AOV $\leq$ 120°
Observation Coverage Ensure valid building visibility within search radius At least one building within 30m
Max. Images per Building Cap the number of images per building $\leq$ 5 images/building
Image-based filters
File Size Remove corrupted or very small images size $\geq$ 20 KB
Semantic Segmentation Check for sufficient visible building surface building_ratio $\geq$ 0.2, wall_ratio $\leq$ 0.3, and vegetation_ratio $\leq$ 0.75
Scene Classification Exclude indoor or irrelevant content environment_type = outdoor
Blurriness Estimate clarity via Laplacian variance blur_score $\leq$ 30
Brightness Filter overly dark or overexposed images 20 $\leq$ brightness $\leq$ 200
4.2 Baselines
Models.
To evaluate the effectiveness of large VLM, we compare it against a set of established CV architectures. These models span both classical CNNs and Transformer-based models to serve as baselines across various tasks. All models are implemented using the PyTorch library and initialized with pretrained weights from ImageNet-1K.
  • VGG: A deep convolutional network characterized by its straightforward architecture of stacked convolutional layers with small receptive fields (Simonyan and Zisserman, 2014). We employ the VGG16 variant in this study, which has previously been applied to building classification tasks (Kang et al., 2018).
  • DenseNet: Featuring dense connectivity between layers, DenseNet facilitates feature reuse and improves gradient flow while reducing the number of parameters compared to traditional CNNs (Huang et al., 2017). We adopt DenseNet201 as a benchmark model, which has demonstrated effectiveness in building material prediction (Ghione et al., 2022).
  • ResNet: We include three variants, ResNet18, ResNet50, and ResNet101, which utilize residual learning through skip connections to enable deeper architectures (He et al., 2016). These models are widely used in building-related recognition tasks (Gouveia et al., 2024; Liang et al., 2024).
  • ViT: Vision Transformers (ViT) divide images into fixed-size patches and use Transformer encoders to capture global context through self-attention mechanisms. Here, we evaluate ViT16, ViT32 and Swin Transformer (Swin_b), which have shown strong performance in previous building classification studies (Raghu et al., 2023; Ogawa et al., 2023).
Each CV model is fine-tuned separately for each building attribute. A systematic grid search is conducted over a range of learning rates (from 1e-7 to 1e-3), and training is performed for up to 64 epochs with early stopping based on validation performance. The checkpoint achieving the best validation metric (accuracy for classification, R-squared for regression) is selected for final evaluation on the test set and used for comparison with the VLM.
Evaluation Metrics.
To comprehensively evaluate model performance across different tasks, we adopt a set of widely used metrics, tailored to the respective output types: categorical labels, numerical values, and text.
For tasks such as predicting building type and surface material, we report four standard classification metrics: Accuracy (Acc), macro Precision (Pre), macro Recall (Rec), and macro F1-score (F1). These metrics have been commonly used in prior works on building classification (He et al., 2024; Sun et al., 2022b; Liang et al., 2024) to evaluate model effectiveness. For continuous attributes such as the number of floors and building age, we use $R^{2}$, Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE). These metrics are standard in evaluating building property predictions (Lei et al., 2024; Wang et al., 2024c). To enable learning from numerical labels, we convert construction year into building age by subtracting the year from 2025 (i.e., age = 2025 – year).
Furthermore, the text generated by ChatGPT-4o on the test set serves as a baseline to evaluate performance improvements in image captioning. In this study, we adopt three commonly used metrics in natural language generation: BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and ROUGE_L (Lin, 2004). BLEU evaluates n-gram precision, while METEOR considers precision, recall, synonym matching, and paraphrase alignment to capture semantic relevance more effectively. ROUGE_L measures the longest common subsequence between generated and reference texts, highlighting fluency and textual overlap. These metrics are widely adopted in remote sensing vision-language tasks (Li et al., 2024a), providing a comprehensive assessment of the accuracy and quality of generated descriptions.
4.3 Generalizability and robustness
To align more closely with real-world building profiling practices, we further conduct comparative assessments involving both VLMs and CV models of their generalizability and robustness. These comparisons target two main objectives: (1) assessing their generalizability to unseen data, and (2) evaluating robustness to heterogeneous noises and degradation.
Generalization to unseen city.
Generalizing CV models to unseen cities remains a significant challenge due to the diverse and unique architectural features across cities (Sun et al., 2022b). VLMs, pretrained on extensive, high‐quality image‐text datasets, demonstrate promising potential to overcome these limitations by leveraging their pre‐acquired semantic reasoning and contextual understanding capabilities. To investigate this potential, we conducted an experiment on building imagery collected from Brussels, comparing the performance of established CNN and ViT architectures against our fine‐tuned VLM. For this evaluation, we retrieve a dataset of 3,687 labeled building images by integrating OSM attributes with buildings detected from Mapillary SVI. The dataset is composed of 3,348 images for building type, 186 for surface material, 1,234 for the number of floors, and 106 for building age. The supplementary material provides the detailed class distribution of this dataset. Due to the limited number of surface material labels across classes, we focus our analysis on the remaining three attributes.
Robustness to varying image quality.
Crowdsourced SVI, unlike standardized remote sensing imagery, frequently exhibit diverse quality issues (Hou and Biljecki, 2022). To systematically evaluate the robustness and stability of the model under such conditions, we adopt the methodology proposed by Hendrycks and Dietterich (2019), which assesses model performance in the presence of common image corruptions and perturbations. Following the image quality criteria defined by Hou and Biljecki (2022), we algorithmically generate four types of corruption to the test set: occlusion, motion blur, Gaussian noise, and brightness alteration, as illustrated in Figure 7. We acknowledge that these distortions are simulated rather than naturally occurring, but they provide a controlled and reproducible way to benchmark model behavior under common visual degradation scenarios.
@@ 4.3 Generalizability and robustness
We then evaluate each model’s performance under these degraded conditions using Relative Corruption Errors ($Relative$ CE) (Hendrycks and Dietterich, 2019). First, the baseline error rate $E^{m}_{clean}$ is determined for model $m$ on the uncorrupted data. Next, we compute the error rate $E^{m}_{c,s}$ for each corruption type $c$ at severity level $s$ (1 $\leq$ s $\leq$ 3). In classification tasks (building type and surface material), the error rate is defined as 1Accuracy1-Accuracy, whereas for numerical values (predicting number of floors and building age), it is defined as $1-R^{2}$. Finally, to account for the varying difficulties introduced by each corruption, we normalize these error rates by dividing by the ResNet50 baseline error. $Relative$ CE is calculated as:
@@
This normalization provides a clearer measure of how much each model’s performance declines under different corruptions. Averaging these $Relative$ CE from four types of corruptions results in the RelativeRelative mmCE, which represents the overall relative performance degradation when the models encountering corruptions.
@@
4.4 Ablation experiments
As one of the very first studies to apply fine-tuned open-source VLMs to the task of building profiling, we conduct a series of ablation experiments to explore and identify effective fine-tuning strategies.
Fine-tuning settings.
To determine optimal training configurations, we first compared two fine-tuning strategies: full fine-tuning, where all model parameters are updated, and parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA). LoRA introduces trainable rank-decomposition matrices into the model while freezing the original weights, enabling more efficient training. This comparison allowed us to evaluate the trade-offs between flexibility, performance, and training cost under different optimization schemes.
We then performed a systematic grid search over learning rates and the number of training epochs. The default learning rate of 4e-5 is used as a baseline, and we evaluat additional values (e.g., 8e-6, 4e-6, and 4e-4) within a predefined range to observe their effects on model stability and performance. We also vary the number of training epochs from 1 to 5 while keeping all other hyperparameters fixed. For each setting, the model is evaluated on a held-out validation set.
Model size.
We compare multiple model variants to assess how model capacity influences performance in the context of building profiling. Specifically, we evaluate InternVL3 models with 1B and 2B parameters, as well as InternVL2.5 model which has 4B parameters avaliable. These experiments are conducted to understand the scalability of different VLM sizes and to determine whether increase model complexity leads to substantial gains across all building profiling tasks.
Data size.
To assess the impact of training data volume, we conduct an ablation study by fine-tuning InternVL3-2B on varying proportions of the complete image-text dataset. We evaluate performance on both attribute prediction and image captioning tasks. This analysis quantifies how model accuracy scales with additional data and offered insights into the marginal benefit of increased data volume in multimodal learning.
5 Results
@@ 5.1 Street-level building dataset
Table 3 provides a detailed breakdown of the number of buildings and SVI images retrieved, individual buildings detected with associated images, and the ratio of completeness for each city. While completeness varies among cities due to differences in the availability and quality of Mapillary images uploaded for specific locations, around 50% of buildings in city centers can be observed and analyzed.
@@
City Total building footprints Total images retrieved Total individual building images Buildings with images Percentage detected City center completeness ($\times2.5km$)
@@
As discussed in Section 3.3, building images with available attributes are sampled to construct a class-sufficient dataset for model development, resulting in a total of 31,180 images. Figure 8 illustrates the distribution of images across categories for each attribute, comprising 17,530 images for building type, 2,871 for surface material, 7,228 for floors, and 5,927 for age.
The dataset is split into training, validation, and test sets in a 6:1:3 ratio. As detailed in Table 4, the training set comprises 20,056 OSM‐sourced labeled samples from 19,443 unique images (some buildings carry multiple attribute labels). Those same 19,443 images are used to generate 19,443 multi‐attribute Q&A pairs and 19,443 descriptive captions via ChatGPT-4o (2024-08-06 release), yielding a total of 58,942 image–text pairs for VLM training. The held-out validation set guides checkpoint selection and hyperparameter tuning for both CV and VLM models based on their performance on multi-attribute prediction tasks. The test set, consisting of 9,016 images with ChatGPT generated Q&A and captions, is reserved for final model evaluation and comparison. To better capture ambiguous architectural features, we also include an auxiliary classification setup by introducing prompts such as “alternate_building_type” and “alternate_surface_material”, which elicit the top-two predictions from the MLLMs.
Table 4: Summary of training, validation, and test data used in this study.
Data Type Source Train Validation Test
Image label
Single-attribute label OSM 20,056 2,737 9,190
Image-text pairs
Single-attribute Q&A OSM 20,056
Multi-attribute Q&A ChatGPT-4o 19,443 9,016
Captioning ChatGPT-4o 19,443 9,016
Total pairs 58,942 18,032
@@
5.2 Model performance